### v1

# Tensor networks for quantum computing

Abstract
In the rapidly evolving field of quantum computing, tensor networks serve as an important tool due to their multifaceted utility. In this paper, we review the diverse applications of tensor networks and show that they are an important instrument for quantum computing. Specifically, we summarize the application of tensor networks in various domains of quantum computing, including simulation of quantum computation, quantum circuit synthesis, quantum error correction, and quantum machine learning. Finally, we provide an outlook on the opportunities and the challenges of the tensor-network techniques.
I Introduction
Field Methods Applications Advantages Challenges
Simulation of quantum computation (Section III) Unstructured TNs, matrix product states (MPS), projected entangled pair states (PEPS), multi-scale entanglement renormalization ansatz (MERA), density matrix renormalization group (DMRG), various renormalization methods Simulation of quantum circuits, analog quantum processors, boson sampling, quantum algorithm benchmarking, quantum many-body physics simulation Simulation complexity reduction, efficiency in representation of some quantum states Representing highly-entangled quantum states, simulating long-time dynamics, simulating deep quantum circuits
Quantum circuit synthesis (QCS) (Section IV) Encoding of TN quantum states into quantum circuits, realizing TN-inspired quantum circuits Preparation of quantum states relevant for quantum computation and quantum simulation Efficiency and interpretability in quantum state preparation Requirement of many-qubit gates whose number normally scales with the bond dimension
Quantum error correction (QEC) (Section V) Tensor-network decoders, TN inspired quantum error-correcting codes Tensor-network decoders of error-correcting codes Efficiency and interpretability in decoding, interpretability in creating and studying error-correcting codes Computational resources overhead, speed, lack of open-source software
Quantum machine learning (QML) (Section VI) Variational tensor-network circuits as quantum neural networks (QNNs) Generative and discriminative models, quantum data classification, quantum data encoding Analytical interpretability, qubit efficiency, noise resilance, absence of barren plateaus, potential reduction in parameters number, avoiding costly TN contraction on classical computer Requirement for nonlocal interactions in quantum hardware for many TN circuits, lack of theoretical guarantees for advantage
Table 1: Summary of tensor network applications in quantum computing.
Tensor networks (TNs) have become a useful tool in many areas of physical and mathematical sciences, especially in the field of quantum information science. The interest in quantum computing (QC) [1, 2] has driven a lot of the development in TNs because they are used to represent and manipulate quantum states and processes.
TNs were initially applied to quantum many-body simulations [3], for which they offer substantial advantages over alternative methods for simulating weakly coupled quantum systems and quantum systems with significant locality [4]. Over time, TNs have broadened their application scope to quantum information theory and quantum chemistry [5, 4, 6], and they have become indispensable to the QC field. In particular, TNs are employed to simulate real quantum experiments that were previously believed to be beyond the capability of classical computers [7, 8, 9, 10]. This demonstrates that TNs successfully address the so-called curse of dimensionality—the problem that the size of the state space increases exponentially with the number of degrees of freedom.
Additionally, TNs provide a general framework of analyzing mathematical objects frequently encountered in quantum information science, which makes them attractive tools to address many other challenges faced by QC. Table 1 provides an overview of TN applications in QC that we will discuss in this paper and spans four subdomains: quantum circuit simulations, quantum circuit synthesis (QCS), quantum error correction (QEC), and quantum machine learning (QML). The table outlines the specific methods used, their primary applications, key advantages, and the challenges they encounter.
We structure the rest of the paper as follows. Section 2 introduces tensor networks, key building blocks, and commonly used methods. Section 3, 4, 5, and 6 discuss the aforementioned domains of application. Finally, we conclude the paper with discussions on the overall advantages of tensor networks, and provide an outlook of the future on how tensor networks may benefit quantum computing.
II Tensor-network methods
Figure 1: Graphical representations of (a) a vector, (b) a 4th order tensor, (c) matrix multiplication, (d) Equation 2, (e) SVD of a tensor, and (f) the definition of isometry.
Tensors are mathematical objects that describe multilinear relationships between other objects. They can be commonly thought of as multidimensional arrays of complex numbers, where the numerical values of the arrays are coefficients describing the relationships. Each index of a tensor indicates mapping to or from an object, and describing the relationship between $m$ objects requires $m$ indices, resulting in a so-called $m$-th order tensor. For example, a vector $(x_{1},\dots,x_{n})$ can be compactly represented in index notation as $x_{i}$ with $i\in(1,\dots,n)$, and is a first-order tensor. A matrix is thus a second-order tensor $x_{i,j}$ with two indices, among other examples.
One can compose the multilinear relationships through tensor contraction. For example, if a tensor $a_{i,j}$ describes the relationship between the objects corresponding to the $i,j$ indices, and a tensor $b_{j,k}$ for those corresponding to the $j,k$ indices, then the resulting tensor describes the relationship between objects corresponding to the $i,k$ indices. This can be described as
$c_{i,k}=\sum_{j}a_{i,j}b_{j,k},$ (1)
where the $j$ index runs over all possible values. We call this contraction of tensors $a$, $b$ at the $j$ index. Matrix multiplication is thus a particular case of contraction between two 2nd-order tensors at one shared index. In general, we could have arbitrary contraction between tensors such as
$d_{i}=\sum_{j,k,l}a_{i,j,k}b_{k,l}c_{j,l}.$ (2)
In this example, a 1st-order tensor $d$ is obtained by contracting a 3rd-order tensor $a$ with two 2nd-order tensors $b$ and $c$.
Tensors and operations upon them are often represented using visual diagrams. An $m$-th order tensor is represented as a node with $m$ edges emerging from the node, each representing an index. Sometimes, the shape and direction of an edge may denote specific properties of the tensor or its indices.
Contraction between two tensors at index $j$ is represented by joining two nodes at the shared edge corresponding to index $j$. Figure 1 provides examples of graphical representations. Since representations of complex contractions between multiple tensors are a network of connected nodes like in Figure 1(d), such mathematical expressions are referred to as Tensor Networks.
Besides composing tensors through contraction, one can also decompose them. This can be done using, for example, the singular value decomposition (SVD). For any matrix $M$, the SVD yields
$M=U\Sigma V^{\dagger},$ (3)
where $\Sigma$ is a diagonal matrix of the singular values and could be absorbed into $U$ or $V$, and $U,V$ are isometries ($U^{\dagger}U=I,V^{\dagger}V=I$). This decomposition presents an opportunity for approximate representation of the original tensor by trimming the singular values. Further, the Eckart–Young theorem [11] states that for approximations with a fixed rank, the solution provided by SVD is optimal. There exist other matrix decompositions which we do not discuss for the sake of brevity [12].
Graphically, the SVD decomposition is shown in Figure 1(e). There, isometries are represented as orangle nodes with directional bonds shown in Figure 1(f). Unitaries, being a special case of isometries, will be represented as magenta nodes with directional bonds as shown in Figure 2.
II.1 Architectures
Figure 2: Different types of TNs: (a) matrix product state, (b) multiscale entanglement renormalization ansatz with the periodic boundaries. The orange circles represent isometric tensors and the magenta circles represent unitary disentanglers. (c, d) projected entangled pair states in 2D and 3D.
The matrix product state (MPS) is the simplest yet the most widely used architecture. In general, the amplitudes of an $N$-body quantum wavefunction form an $N$th-order tensor, and the MPS provides an approximate representation of this tensor as
$c_{i_{1},\dots,i_{N}}=\sum_{\alpha_{1},\dots,\alpha_{N}=1}^{\chi}\Gamma_{\alpha 1}^{[1]i_{1}}\Gamma_{\alpha_{1}\alpha 2}^{[2]i_{2}}\dots\Gamma_{\alpha_{N-1}}^{[N]i_{N}},$ (4)
which is a contraction of the $\Gamma_{i}$ tensors, each corresponding to a site in the system as shown in Fig. 2 (a). Here, $\chi$ is the parameter called the bond dimension which controls the accuracy of compression. This TN is especially effective for simulating one-dimensional (1D) quantum systems. It follows the area law of entanglement [13], but has been used for other systems as well [14, 9]. Its contraction cost is polynomial in the system size, meaning the physical observables can be evaluated efficiently. This concept has independently emerged in computational mathematics, where it is called the tensor train [15].
Projected entangled pair states (PEPS), shown in Fig. 2 (c) and Fig. 2 (d), aim at generalizing the MPS to higher dimensions, making them capable of reproducing large entanglement and suitable for simulating high dimensional quantum systems. However, algorithms for PEPS are typically much more computationally demanding [5, 16].
The multiscale entanglement renormalizaton ansatz (MERA), shown in Fig. 2 (b), is designed to capture the entanglement structure of quantum systems across different length scales and can reproduce logarithmic violation of the area law.
The above TNs can also be equipped with the periodic boundary conditions. Furthermore, they can be generalized for providing the description of operators such as Hamiltonians and quantum channels. One can even define TNs with arbitrary geometries for special applications. For example, the quantum state of a quantum circuit with single- and two-qubit gates can be represented by a network reflecting the connectivity [17]. Consequently, the circuits with gates laid out in a fixed geometric pattern (e.g., a 2D grid) may be easier to simulate [17, 10, 18].
II.2 Manipulation of TNs
While TNs can be used to represent a broad range of mathematical objects in quantum information science, their popularity stems primarily from their power as a tool for obtaining meaningful, interpretable quantities such as operator expectation values, reduced density matrices, amplitudes, samples, and more. Due to the high interest in the TN techniques, a plethora of software libraries suitable for different types of TN manipulations are available, such as cuQuantum [19], Cotengra [20], iTensor [21], and many more [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
II.2.1 Update
One of the important tasks is optimizing a TN with respect to a cost function, such as the energy of a quantum state. A pioneering approach to this problem was introduced by Steven White in 1992 in which the density matrix renormalization group (DMRG) [37] is used. For a cost function defined by a Hamiltonian on a lattice, DMRG is a variational optimization technique that finds the best MPS approximation of the many-body wavefunction of the ground state. This algorithm performs global optimization by sequentially optimizing each local tensor (or pairs of them, depending on the particular version of the algorithm), which is repeated across several sweeps.
In addition to the DMRG algorithm, gradient-based optimization methods can also be employed if the gradient of the objective function is accessible, for instance, through automatic differentiation [38, 39]. Additionally, Riemannian optimization [40, 41] can be particularly helpful when one needs to maintain the isometric properties of the TN [42, 43].
Another class of methods relies on Monte Carlo sampling. In the variational Monte Carlo (VMC) approach, observables are evaluated by sampling configurations of the many body system instead of exact calculations [44, 45].
One may also consider the time evolution of a quantum state or an operator in the TN representation. For discrete time steps, this can be done by applying local gates or Kraus operators through contraction, and restore the original form of the TN with SVD. These approaches are called time-evolving block decimation (TEBD). Continuous time evolution under a Hamiltonian can be approximated by discrete Trotter steps and the discrete method can be applied if the Hamiltonian is sufficiently short ranged. For long-range Hamiltonians, TEBD becomes inefficient, and methods such as MPO $W^{II}$, TDVP, and Krylov approaches are preferred; see [46] for a recent review. Krylov methods use Lanczos diagonalization to compute the action of the time evolution operator, while TDVP evolves the state directly on the MPS manifold for fixed bond dimension [47, 48, 49, 50].
II.2.2 Contraction
A crucial part of many TN algorithms is the contraction – the evaluation of a single scalar or tensor represented by the network. Naively, the cost of the contraction scales exponentially in the number of indices. In practice, the contraction can be performed by using a sequence of intermediate tensors using pairwise contractions, known as an ordering or a contraction tree. The optimal contraction cost of this approach still generally scales exponentially [17], but the approach nonetheless offers a dramatic reduction in the cost at the expense of some intermediate memory. Notably, certain tree- and fractal-like geometries can be contracted with only polynomial cost [51].
Finding the optimal contraction scheme is itself an NP-complete problem [52, 53]. However, much progress has recently been made using heuristic approaches targeting the total cost of the contraction operation. Specifically, these methods have been based on recursive graph partitioning [54, 20, 55], simulated annealing [56, 57] and now reinforcement learning [58, 59]. The extraordinary sensitivity of the cost to the contraction tree quality has led to improvements by many orders of magnitude for some problems such as, for example, quantum circuit simulation.
For exact simulation of quantum circuits, two additional techniques are central to achieving a state-of-the-art performance. First, whilst the time cost of a contraction might be acceptable, the space cost in terms of intermediate memory might still be enormous. Slicing (also known as cutting or projecting) splits a contraction into many smaller independent contractions [60, 61, 62, 63, 64], each of which for example can fit on a GPU and be performed in a massively parallel way [65, 56]. Depending on the geometry, this can sometimes be performed with very little overhead. Secondly, if one wants to evaluate many related tensor networks differing only in some entries (for example, a set of amplitudes), one can employ multi-contraction [66, 7] to cleverly avoid repeating the same computation. The largest exact simulations of quantum circuits use both of these techniques in tandem [7, 67, 68, 69].
To go beyond exactly treatable network sizes, one must use approximate contraction. While even this is not expected to be viable in the general case [70], evidence obtained from simulating many-body physics systems shows that many real world tensor networks are tractable approximately [71, 72, 73, 74, 75, 76]. Time evolved MPS [77, 50] and PEPS [78, 79, 80, 81, 10] can be thought of as approximate contraction, in which case one is limited by buildup of entanglement. These methods have historically been handcrafted, but recent work has focused on automatic approximation [82, 6, 83] and contraction sequence-optimized [84] approximate contraction. An outstanding question is in which classes of quantum circuits do such techniques admit a polynomial or exponential cost reduction [85].
III Simulation of quantum computation
Many quantum objects—statevectors, operators, channels, and others—can be represented as tensors. Given that quantities of interest often derive from these objects, TNs are often useful for simulating quantum systems.
III.1 Gate-based quantum computation
TNs can be used for simulating quantum algorithms in the gate-based model. One of the approaches includes evolving an ansatz, such as an MPS or PEPS, using TEBD-like techniques. Alternatively, a quantum state can be represented as a TN of contracted circuit gates with fixed input indices and open output indices. By contracting appropriate TNs constructed from the state network, amplitudes and expectation values can be computed. Another example is the evaluation of the trace of quantum circuits, which enables the estimation of the circuit ensemble randomness [86].
A good example is the simulation of quantum circuits used to simulate many-body physics, such as the experiment on the dynamics of a kicked Ising model [87]. While the problem was considered initially intractable using state-vector MPS and isometric-TN [88] approaches, it was later demonstrated that TN belief propagation [10], Heisenberg MPO [89] and PEPO time-evolution [90], gPEPS [91], as well as noisy [92] and sparse Pauli dynamics [93] methods can solve the same problem more efficiently, and can even produce more accurate results than the quantum processor itself.
Another important class of experiments suitable for TN techniques is a random circuit sampling (RCS). RCS is a computational task where a quantum computer executes a random $n$-qubit quantum circuit and output the measurement results (a length-$n$ bitstring) of the resulting quantum state. This task is classically hard because the existence of an efficient classical algorithm for estimating the probability of a given output would lead to the collapse of the polynomial hierarchy [94, 95]. As a result, RCS has been experimentally demonstrated many times as a proof of quantum computational advantage [96, 97, 98, 57, 18]. Additionally, applications based on RCS in which classical simulation using TNs consistitute an integral part have been proposed [99, 100].
The leading approach for simulating RCS with the lowest computational cost and the highest fidelity is contracting the TN representing the quantum circuit as a network of contracted tensors that correspond to the applied gates [17]. Contracting a fraction of slices effectively performs a finite fidelity simulation [56]. Approximation and simplification of the original TN could also be implemented alongside slicing, which reduces the fidelity [67]. Additionally, conventional RCS experiments usually request millions of samples from a single circuit. One can reduce the cost by leaving some indices open [101, 102, 7], using a sparse output state [67], or reuse intermediate results [56, 66, 7]. One can also perform some post processing on the samples to spoof the quality metric [69].
Another commonly analyzed approach is the MPS based on DMRG [103]. This approach also breaks the TN corresponding to a sample amplitude into three parts: the beginning of the circuit with fixed input indices, the middle, and the end with fixed output indices. The first and third parts can be represented using an MPS, and the amplitude can be obtained by contracting the two MPSs with the middle part in between.
Finally, prior to the development of the aforementioned methods, other TN techniques such as MPO [104], MPDO [105], PEPS [106], conversion of 3D to 2D networks [63] are utilized.
III.2 Analog evolution
Analog quantum computers are devices where a specific class of models is implemented “natively” on the quantum hardware. These systems include optical lattices of neutral toms [107], trapped ions [108], Rydberg tweezers [109], photonic waveguides [110], as well as superconducting circuits [111]. Contrary to digital quantum processors, analog quantum devices allow the simultaneous, time-dependent continuous control of pairwise interactions of all the qubits available on the quantum chip. The major drawbacks of analog quantum simulators are the calibration errors of Hamiltonian parameters and decoherence. One also needs to make sure that the quantum processor is faithfully behaving according to closed-system Schrödinger time-evolution for the largest number of qubits possible. TNs provide an invaluable benchmarking tool in this regard.
In adiabatic quantum annealing [112] (AQA), the lowest-energy state of a complex Hamiltonian is sought by starting from a simple and well-defined initial low-energy state of a well-controlled Hamiltonian. Then the parameters of the established Hamiltonian are changed very slowly to arrive at the more complex Hamiltonian. In a coherent QA, the annealing process is performed faster, i.e., on the time scale well less than the expected qubit decoherence time. Therefore, rather than seeking for optimization of the final complex Hamiltonian, one simulates the dynamics of a closed quantum system swept through a quantum critical point. In particular, Ref. [113] claimed that in the studied the parametric range, the approximate classical TN methods such as MPS and PEPS cannot match the solution quality of the quantum simulator in solving the Schrödinger dynamics for a transverse field Ising spin glass system in 3D and biclique (all-to-all) lattice geometries, despite the limited correlation length and finite experimental precision. Time evolution of the MPS was performed using the GPU accelerated TDVP on snake-like unfolded 2D, 3D, and biclique lattices (which was found superior against other methods, such as TEBD as well as local and global Krylov methods [46]) performed on Summit and Frontier ORNL supercomputers. In this study, the MPS methods played a crucial role via the estimation of an equivalent “QPU bond dimension” [114], defined by matching the sampled QPU distribution quality against converged MPS simulations at simulatable scales.
A recent simulation technique was able to achieve comparable accuracies to the quantum annealer [115] for two- and three-dimensional systems. It uses belief propagation for time evolution, and more sophisticated variants of belief propagation for calculating expectation values. However, this method is not immediately appicable to the inifite dimensional biclique lattice studied by the quantum annealer.
III.3 Boson sampling
Boson sampling [116] is a computational model based on linear optical elements, non-linear input, and measurements. This model of computation is non-universal, but it is, nonetheless, hard to simulate classically under plausible complexity theoretic assumptions. Similar to RCS, boson sampling is a computational task of producing samples from a probability distribution corresponding to measurements of the outputs of a linear-optical interferometer.
Since the transmission matrices of boson sampling experiments are approximately Haar-random [117, 118, 119, 120], using an MPS approach is suitable despite the high dimensionality of the optical interferometer [121, 122]. Additionally, for simulations without photon loss, photon number conservation implies a $U(1)$ symmetry, which is further exploited to reduce the cost of the MPS [121, 122]. In reality, various experimental noises, especially photon loss, scale with the system size, and the complexity-theoretic argument is not applicable in this case. Therefore, directly approximating the mixed state in the Fock basis using an MPO can be efficient. In this approach, if we have $N$ input photons and $N_{\rm out}\propto N^{\gamma}$ output photons (where $\gamma$ is the scaling exponent), the MPO entanglement entropy (MPO EE) grows as $S=O(N^{2\gamma-1}\log N)$, which indicates a logarithmic growth of the MPO EE when $N_{\rm out}\propto\sqrt{N}$[122, 123].
Further, for Guassian boson sampling [124], the lossy output states can be modeled by applying random classical operations on a pure state [9, 125]. This allows the simulation of the pure state with an MPS. Further, one can optimize this pure state to have significantly fewer photons compared to the final output state [9]. The bond dimension grows logarithmically when $N_{\rm out}\propto\sqrt{N}$, and this theoretically guarantees polynomial growth of the bond dimension to fixed fidelity, allowing the largest boson sampling experiments to be simulated on up to $288$ GPUs in under two hours [9].
Additionally, one can simulate boson sampling in the Heisenberg picture in a way similar to the simulation of quantum circuits. Instead of evolving the quantum state using a TN, this approach evolves the Fock state projector of the desired measurement outcome [126]. For Gaussian boson sampling, the Schrödinger picture approach evolves a Gaussian state, which usually has a larger bond dimension, whereas the Heisenberg picture evolves a Fock state.
IV Quantum circuit synthesis
Quantum circuit synthesis is the process of decomposing a target quantum operation into a sequence of executable gate operations that are compatible with a specific quantum computing architecture [127]. When considering modern quantum processors, the problem of quantum circuit synthesis faces two challenges: (1) the decomposition algorithms must adhere to the native connectivity of the quantum device and (2) the circuit depth allowed to faithfully prepare a quantum operation is limited by the characteristic noise in the quantum device. A universal solution of these problems, either in terms of scalability or precision, remains, at the present stage, out of reach.
TNs offer a promising general pathway for addressing these challenges [128]. TN architectures, such as MPS and MPO, feature an inherently geometrical layout provided by the prescribed decomposition/representation of general tensor formats. It turns out that for circuit descriptions, which inherently are TNs, such layouts are particularly attractive due to the ease in attaining connectivity within adjacent sets of qubits. Although this feature is exploited in studying long/short range interactions in condensed matter systems, it naturally provides a partial solution to the problem of circuit synthesis in terms of realizing native hardware connectivity. By representing quantum states and operations as TNs and further casting them as circuits, one can simplify both circuit design as well as address the problem of compiling arbitrary unitaries into natively realizable gates. This can reduce the complexity of quantum circuits and enable more efficient synthesis.
IV.1 Promoting TNs to quantum gates
Quantum states and channels, represented as TNs can be mapped to quantum circuits. The process of promoting TNs to quantum gates requires four steps: (i) transforming the original TN into a TN of isometric tensors; (ii) embedding spatial and temporal directions to the network; (iii) promoting each isometric tensor into a unitary; and (iv) decomposing each unitary as quantum gates.
Quantum operations comprise unitary operations across multiple qubits. In order to map a TN to a quantum circuit, each tensor in the network is mapped to a unitary tensor. An intermediate step in this approach is to convert every tensor in the network into an isometric tensor. Ensuring that each tensor is an isometry is possible due to the gauge freedom in TNs [129, 130]. Additionally, quantum circuits are inherently directional in time, whereas TNs have no directionality. Therefore, we are free to translate and rotate each tensor in the diagram. In order to turn TN diagrams into QC diagrams, we need to place an arrow of time onto the TN diagram, and specify incoming and outgoing wires. Once a TN comprises only isometric tensors, we promote the individual tensors into unitaries. This procedure is depicted in Fig. 3(a).
IV.2 Examples of a TN state and operator preparation
Using the introduced prescription, we provide examples of important TN archectures in the literature, and demonstrate how they can be mapped to quantum circuits. Figure 3 depicts how one converts an MPS into a quantum circuit using the steps from the above subsection. The key element in this procedure is that an MPS can be exactly transformed into a canonical form, including only isometric tensors, which is a unique feature of an MPS—this is not the case for the higher-dimensional TNs like PEPS. Preparing an MPS as a quantum circuit was first introduced in [131], and has since been theoretically explored by many others [132, 133, 134, 135, 136]. An experimental realization of an MPS has been extensively explored across various quantum computer architectures [137, 138, 139, 140].
Figure 3: (a) Promoting an isometric $d_{1}\times d_{2}$ matrix $V$ into a unitary gate $U_{V}$, with an incoming $p$-dimensional ancilla qudit and an outgoing $q$-dimensional qudit that is measured. The dimensions of $p$ and $q$ are $p=\text{lcm}(d_{1},d_{2})/d_{1}$ and $q=\text{lcm}(d_{1},d_{2})/d_{2}$, where lcm denotes the least common multiple operation. (b) Same as (a), but the isometry is taken from an MPS tensor. (c) Converting a (canonized) MPS into a quantum circuit by promoting isometric tensors into unitaries. Each individual unitary is then decomposed into one and two-qubit gates by encoding the qudit wires to qubit wires followed by methods such as Gray codes to decompose arbitrary unitary matrices into two-level unitaries. The arrows in the figure represent the isometric conditions of the tensors, as defined in the convention in Figure 1, and do not indicate circuit time.
Although such methods are accurate, an exact MPS preparation requires unitary operations acting immediately on the $(\lfloor\log(m)\rfloor+1)$ qubits, where $m$ is an MPS bond dimension. This may be undesirable for modern quantum computers, given their limited connectivity and the restricted set of native hardware gates. One approach to address this issue is to variationally fit a network of local circuits composed of native gates in order to approximate the original unitary tensor network. Alternatively, iterative methods can be used to remove short-range entanglement through native-gate disentanglers [138, 141, 136, 133].
The PEPS, however, generally lack a canonical exact isometric form due to the higher connectivity inherent in their underlying structure. In other words, one cannot exactly map a PEPS onto a quantum circuit without an exponential amount of postselection. However, one can create a subclass of PEPS, referred to as isoPEPS, to establish a proper connection to quantum circuits. In the isoPEPS class, each tensor, or set of tensors, respects the isometric condition, making it equivalent to a quantum circuit [88, 142, 143]. The question of undestanding the varational power of isoPEPS and the way how it is compared to PEPS and to 2dMERA and how they can be realized experimentally on a digital quantum computer is now an active area of research.
Among all tensor network methods for classical simulation of quantum systems, the MERA stands out as the most natural representation of a quantum circuit: it was originally envisioned in the reverse direction of a quantum circuit in which all isometries are embedded in unitaries. It features an intrinsic robustness to noise and does not suffer from barren plateaus [144, 145, 146]. The MERA has recently been demonstrated on the ion-trapped digital quantum computer to probe the criticality of many-body systems [147, 148, 149].
IV.3 Implementation techniques
As mentioned in Sec. 4.1, TNs cannot be directionality compared to real quantum circuits. Because of this, we must choose how tensors are executed in time, which then is related to the quantum circuit architecture. This freedom allows TN states to be prepared holographically. Introduced first in [137], a holographic preparation of a TN state takes a spatial dimension of the physical state and prepares it sequentially in time, with the important feature that once a circuit is done preparing one site, the qubit can be reused to prepare the next site.
In addition to the holographic implementation techniques where each site is sequentially prepared after the next site, adaptive circuits have recently been explored to prepare TN states [150, 151, 152, 153, 154, 155, 156]. Adaptive circuits are quantum circuits that allow mid-circuit measurements followed by gate operations that are determined by those measurements. While pure unitary circuits require a circuit depth scaling with system size to prepare states with long-range correlations, adaptive circuits can prepare certain states in constant-depth.
Beyond the exact implementations of mapping a TN to a quantum circuit described above, various variational methods are employed to prepare an approximate TN state [157, 158, 159]. The basic idea is to numerically optimize a variational ansatz to approximate the target state, typically resulting in a shallower quantum circuit. In addition to preparing a TN state, the work [160] used variaional methods to instantiate a generic quantum circuit where the TN formulation is leveraged as an efficient back-end in the compilation workflows.
V Quantum error correction and mitigation
Quantum error correction (QEC) plays a critical role in safeguarding quantum information from errors caused by decoherence, dissipation, and control inaccuracies [127]. In this section, we discuss two main topics: (i) the parallels between quantum error-correcting codes and TNs, along with (ii) the application of TNs in decoding, including the formal reduction of optimal decoding to a TN contraction [161].
V.1 Tensor-network codes
The first major intersection between TNs and QEC was the establishment of a formal correspondence between certain QEC codes and the associated TNs [161]. Convolutional codes, concatenated block codes, and topological codes, for example, can be respectively represented by the MPS, tree TNs, and PEPS. This insight allows for importing well-established tools of TNs to represent and analyze QEC codes. In particular, the problem of an optimal decoding of an QEC code (i.e. diagnosing errors, see the discussion in Section V.2) was shown to be formally equivalent to contracting an associated TN [161] approach. For convolutional and contatenated block codes, which have efficiently contractable TNs, this correspondence thereby provides the efficient optimal decoders.
The QEC-TN correspondence was later generalized by the introduction of TN stabilizer codes [162] and the ‘quantum lego’ formalism [163, 164, 165] for constructing large codes out of a finite set of smaller ‘seed’ codes. The seed codes are representend by small tensors that are combined into a TN to represent a larger code. This construction was shown to be universal in the sense that any qudit QEC code can be represented by a TN built out of three elementary seed tensors [163]. As before, TN codes come naturally equipped with optimal decoders that are evaluated by contracting a TN [162].
As an important special case, TNs provide a powerful formalism for representing arbitrary topological and holographic codes [162, 166]. For example, Ref. [166] showed how simple modifications of the TN for a topological code can provide codes with more favorable properties. TN codes also provide a more general framework for better understanding holographic QEC codes such as those constructed out of isometric [167] and hyperinvariant [168] tensors that are featured in models of quantum gravity and holography, and can additionally exhibit favorable properties for practical QEC [169].
V.2 Syndrome decoding
Decoding a quantum error correcting code involves processing syndrome measurements using a given theoretical error model. This process yields appropriate correction operations and involves solving the inference problem of finding the set of errors that ocurred in the computation with the highest likelihood based on the syndrome information measured and on the error model at hand. This process has been shown to be formally equivalent to contracting a TN [161, 170].
The error model of a quantum error correcting code provides two sets of information: 1) the probability of occurrence of the different physical errors and 2) the recognition of what syndrome measurements each of these errors affects and how each one affects the logical qubit(s) encoded. Given a set of syndrome measurements, there are exponentially many configurations of physical errors compatible with it. This set of error configurations can be further decomposed into several subsets, defined after their effect on the logical information encoded; these subsets are often called cosets. The sum of the joint probabilities of the error configurations of each coset is proportional to the likelihood of the coset. Optimally, a decoder infers the maximum likelihood coset as the most probable logical outcome of the code, given a set of syndrome measurements. This is an exponentially large sum of probabilities computed, in a good approximation, by the contraction of a TN.
There are two approaches to carry out this summation. The first and most popular one consists of identifying a complete set of symmetries of the coset, i.e., all sets of errors that leave the logical information and the syndrome mesasurements unchanged. In this setting, an enumeration of all error configurations included in the coset is achieved by finding a particular configuration in the set and applying to it all combinations of the symmetries [161, 170, 6, 171, 172].
A second approach consists of building a TN directly from the error model connectivity between errors and syndrome detectors. In this approach all error configurations outside a coset are not considered by explicitly zeroing out tensor entries corresponding to them. This approach has the benefit of working with more general error models directly, without the need to find a set of symmetries that is computationally beneficial [173, 174, 175].
Optimized decoding algorithms benefit from the computational strategies inherent in TN theory, such as the use of approximate contraction methods for large-scale networks, which are directly applicable to decoding large and complex QEC codes [6, 171, 172, 174, 173, 175]. TN decoders are, up to date, considered as highly accurate but not practical for real time decoding [176]. The need for fast decoders in real quantum error correction experiments makes the use of heuristic decoders optimized for performance more popular and practical. This leaves TN decoders as useful tools for the benchmark of experiments and other less accurate decoders.
V.3 Quantum error mitigation
Error mitigation methods are expected to play a crucial role in near-term quantum experiments until fault-tolerance is fully achieved [177, 178]. One of the promising methods in this area is probabilistic error cancellation (PEC). However, it requires precise knowledge of the underlying noise model at the gate level. PEC creates an ensemble of circuits, with a denoiser, which on average replicates a noise-inverting channel, inserted according to a quasi-probability distribution, effectively canceling out the noise. A quasi-probability distribution is a probabilistic framework that assigns a sign to each sample, allowing for the representation of non-physical channels. The number of samples (measurement overhead) required to address sign cancellations can be effectively bounded by Hoeffding’s inequality, growing exponentially with the number of qubits and circuit depths. Despite the overhead, it is anticipated that error mitigation will remain practical for circuits with qubit counts in the hundreds and equivalent circuit depth [87]. The tensor-network error mitigation (TEM) algorithms represent a similar paradigm for reducing overhead. As investigated in a series of recent works [179, 180, 181, 182], TEM provides a universal lower cost bound for error mitigation, specifically a quadratic cost reduction. TEM is a post-processing procedure that acts on a set of randomized local measurements, as designed to cancel errors by incorporating classically noise-inverting channels. The mitigiated estimations are given by contracting the circuit-level TN. For systems containing more than a thousand qubits with equivalent circuit depth, we foresee the development of hybrid approaches combining quantum error correction with error mitigation techniques [183]. It remains an open question whether error mitigation alone, or in combination with quantum error correction, can lead to utility in near-term quantum computation or provide a potential practical advantage in quantum algorithms.
VI Tensor networks for quantum machine learning
Over the past decade, TNs have gained significant interest in machine learning (ML). By compressing high-dimensional linear layers in neural networks, TNs reduce memory usage and the number of training parameters [184, 185, 186]. Moreover, TNs offer strong analytical interpretability and are related to various established ML techniques, such as Boltzmann machines [187, 188], Born machines [189], autoregressive models [190], and hidden Markov models [191]. The architectural flexibility of TNs allows the introduction of inductive biases into ML models based on intrinsic data topology and correlations [186, 192, 193, 194], ultimately enhancing training efficiency and model generalization. This synergy has paved the way for employing TNs across various machine-learning applications, including discriminative, generative, and regression frameworks.
Although the contraction complexity of many TNs grows polynomially with the bond dimension, the degree of the polynomials is higher than three in some cases (such as for MERA), which is often regarded as inefficient in practical ML. Quantum computers may assist TN-based ML by implementing parameterized quantum circuits (PQC) reflecting specific TN architectures, resulting in TN-based quantum ML [195] (TN-QML) which avoids expensive contractions. Moreover, in the context of quantum simulation, parameterizing TNs with quantum gates has been shown to offer an exponential reduction in the number of parameters required to achieve results similar to those of classical TNs [138, 141], which could potentially translate to ML tasks as well. On the other hand, such variational TN circuits offer advantages over commonly used quantum hardware-efficient ansätze, such as being free from barren plateaus [196, 197, 145] (flat regions in the optimization landscape that impede training). Additionally, the TN structure yields qubit-efficient implementations [198] and robustness against decoherence [198, 199, 200].
Consequently, quantum versions of MPS [198, 201], TTN [198, 199, 202], and MERA [199, 202] have been explored in QML for both discriminative [198] and generative [201] tasks. The TN-QML approach also inherently enables quantum applications, such as quantum phase classification directly on quantum devices using quantum convolutional neural networks (QCNNs) based on the TTN and MERA architectures [203, 202]. However, although the QCNN was initially considered a candidate for exponential quantum advantage, recent work [204] has argued against this, suggesting it can be dequantized using classical shadow techniques.
However, TNs are promising in QML as a tool for encoding classical data into quantum circuits [205, 206] and for pre-training the TN-PQC on classical computers [198, 207, 208, 209] before extending to quantum ansätze that cannot be simulated classically. Finally, TN surrogate modeling can be employed to test QML moddels for dequantization [210]. For an in-depth review of using TNs in the QML, we refer the reader to [195].
VII Discussion and Outlook
TNs are a valuable tool for the field of quantum computing. They can be important for analyzing computational complexity classes. For example, finding a polynomial-time tensor network algorithms for a problem [211] proves that the problems is in the complexity class $\mathcal{P}$. Similarly, one can empirically argue that a problem is hard by trying existing state-of-the-art TN techniques and observing exponential scaling. Along the same vein, TNs can help validating quantum advantage claims of finite-size quantum experiments, which may not admit asymptotic quantum advantage. Here, rigorous benchmarking of TN techniques as well as other classical techniques allows to decide whether a quantum experiment is out of reach for classical computers.
TNs have proved their value in quantum algorithm analysis, circuit synthesis, and quantum error correction and mitigation, allowing them to enhance the capabilities of quantum computation. For quantum algorithm analysis, TNs are often used to simulate a quantum algorithm and extrapolating the performance scaling. For quantum circuit synthesis, TNs allow for certain states or operators to be prepared on hardware efficiently with limited connectivity. For quantum error correction, TNs are valuable theoretical tools enabling the construction of certain families of codes. TNs can also be used for constructing optimal decoders. For quantum error mitigation, TNs enable an overhead reduction. In QML, TNs are used to design data-specific quantum models that improve resource efficiency, trainability, and robustness against noise.
TNs continue to drive progress across multiple areas of quantum computing as demonstrated throughout this review. As we transition towards fault-tolerant quantum devices, TNs will likely maintain their significance, particularly in potential hybrid quantum-classical devices which are anticipated to integrate quantum computers with specialized classical accelerators like the GPUs, TPUs, and FPGAs.
Looking forward, we foresee that TNs will evolve to include new algorithmic approaches and techniques to meet the demands of fault-tolerant quantum computers. As quantum hardware capabilities improve, the synergy between TNs and quantum computing may deepen, keeping TNs an important tool for further develoment of future quantum computing systems.
Acknowledgments
A. B., R. E., V. K., C. M., A. M., F. N., D. M., M. Pere., M. Pf., and V. V. were supported by Terra Quantum AG. A. N. was supported by Natural Sciences and Engineering Research Council of Canada (NSERC) Alliance Quantum Program (Grant ALLRP-578555), and the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program. S.Ṁ. is partially supported by the Prime Contract No. 80ARC020D0010 with the NASA Ames Research Center and acknowledges funding from DARPA under IAA 8839. A. B. thanks Ilia Luchnikov, Mikhail Litvinov and Stefanos Kourtis for valuable discussions. A.A., Z.H., M.L., M.A.P., R.S. and M.P. thank their colleagues at the Global Technology Applied Research center of JPMorganChase for support and helpful discussions.